11) find the point of diminishing returns of the function 2x^3 - 1.5x^2 +3x
Answer #1

f(x)= 2x^3-1.5x^2+3x

To find the point of diminishing returns, set the second derivative equal to zero.

f'(x)= 6x^2-3x+3

f"(x)= 12x-3=0

x=3/12=0.25

Point of diminishing returns is x=0.25